Session: Sep 1997 (session 3)
Frequency: 1.3 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Tugnoli Marco at Fri Sep 12 08:05:27 METDST 1997) Success

Cambridge (Alastair Gunn at Fri Sep 12 10:11:47 METDST 1997) Minor failures
Vacuum lost, missed pass 1F on tape 2, re-positioned for pass 2R. Tapes jpl05235,0c85 and asi00002,a028. No other known problems. Weather at Cambridge cloudy and wet.

Onsala (Antonis Polatidis at Fri Sep 12 17:41:51 METDST 1997) Success
No known problems. Light rain during some parts of the experiment.

Effelsberg (dgraham at Mon Sep 15 10:52:36 METDST 1997) Success

Metsahovi (Kirsi Karlamaa \& Kaj Wiik at Tue Sep 16 13:30:52 METDST 1997) Success
Tape head might have been in wrong position during 18:00 - 18:06. Otherwise successfull.

Noto (C.Trigilio at Mon Sep 22 16:31:55 METDST 1997) Success
No problem reported

Calibration (Fredrik T Rantakyro at Wed Nov 5 15:42:37 MET 1997) Minor failures
The ANTAB formatted TSYS information can be found on VLBEER in the SEP97 directory. The file is called ed011.antab and contains the latest antenna calibration information.